Let X be exponentially distributed with parameter 3.

a) Compute P(X > 6 | X > 2).

b) Compute E(7e-12x+8+ 5).

c) Let Y be independent from X. Suppose the PDF for Y is f(x) = 2x for 0 ≤ x ≤ 1 (and 0 else). Find the PDF of X + Y.
Answer #1

- Page 11 I Page 1 Cheag ga). Xu Exp (3). ☆ tx) of 3e 3x aro Yo, otheodorce. : PEx>6/872] =P [X>60x72] P[x2]. =P[x>6] P[X72].I PAGE 2 I tehasot Page 2 -stit = 36 ) [230] -= 0-72-]= xe-a-1= (x2 Cara-1-1=[22x19LXd Gisa Del e - 6 . =é 18+6– 12 – 6.14421PASES I Page 3 Cheaal $ ca P[x76\x72J = 6.144212353x10-6 =0.000006144212 353 So I Le 12*+*+5]. E le 12te] +5. - Fietus blandi

PASES I Page 3 Cheaal $ ca P[x76\x72J = 6.144212353x10-6 =0.000006144212 353 So I Le 12*+*+5]. E le 12te] +5. - Fietus blandiT Page 4 cheap I Page 41 8b). -1578 dx +5 + S -70[2+1+5 = 700 [°-17 to. - see +5. 2798 +5 + BC = 4178.34182I Page 5 Page 5 - Cheap 9002 E le 2x+8+5] 54178.341182. Let us de tine U = aty. osy. . » na uro yao Noot (4,0) blacy) lJ. whePage 6 - Cheap I T Page 6] . 6 (4,9)= 3е зи-) 2 Поо аmе o6 49- с. .: + 4) = (+) до 3r dur - Зее 2 | e = Be-34 Г. re — ве за ГI PAGE 7 chega] I Page 1 och, = 6e-29 Pēs (44)– (+)] = 6e-34 [es lot] = Ben Bu (102-1). t(x+4)= (14)= 2e 34 (10031) 3 3 for l